40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 1957     Entry time: Wed Jul 1 17:31:16 2015
Author: Arjun 
Type: Misc 
Category: PD noise 
Subject: Updates on PD noise setup 

In the last few days we were able to get almost all the optics set and aligned. The following few images show the complete setup with all optics aligned and a image showing the laser beam entering the box.

The experimental setup with all the mirrors and optics aligned
The experimental setup outside the box

                                                                                                                                                                               

 

 

 

 

 

 

 

 

 

 

The laser beam entering the box

The next thing to do would be to setup the photodiodes and the readout electronics. Since the digital system the lab is not up yet we decided to hardwire the required filters on the LIGO- generic filter board, becasue using SRS560's was making the space too clumsy.

Attachment 3: IMG_20150630_192049.jpg  1.578 MB  | Hide | Hide all
IMG_20150630_192049.jpg
ELOG V3.1.3-